# Normalized solutions to mixed dispersion nonlinear Schrödinger system with coupled nonlinearity

### v1

@@ 1 Introduction
problem (1.1)-(1.2) has a mountain pass type nontrivial radial solution for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
Here we assume $N\geq 2$, because in Section 4, we will replace $H^{2}(\mathbb{R}^{N})$ by the radial subspace
$H_{r}^{2}(\mathbb{R}^{N}):=\{u\in H^{2}(\mathbb{R}^{N}):u\mbox{ is radially symmetric}\}.$
The compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ for $2<p<2^{**}$ only hold when $N\geq 2$.
This paper is organized as follows. In Section 2, we present some useful lemmas. In Section 3, we give the proof of Theorems 1.1 and 1.3. In Section 4, we give the proof of Theorem 1.5. In the rest of this paper, unless otherwise specified, we always assume that $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $r_{1}$, $r_{2}>1$.
@@ 2 Preliminaries
In this section, we are devoted to some notations and preliminary results.
@@ 3 Mass-subcritical and mass-critical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r=r_{1}+r_{2}\in(2,\bar{r}]$. Initially, we obtain some important properties about $m(\rho)$ given in (1.7).
@@
Let $m^{J}(\rho)$ be given in (1.8), which is well-defined for all $\rho>0$. The following result can give the accurate value of $m^{J}(\rho)$, and it is not attainable.
@@
$-\frac{\alpha_{1}^{2}}{8}\rho^{2}=J(u,v)\geq M(\rho_{1},\rho_{2})=-\frac{\alpha_{1}^{2}}{8}\rho_{1}^{2}-\frac{\alpha_{2}^{2}}{8}\rho_{2}^{2}\geq-\frac{\alpha_{1}^{2}}{8}\rho^{2},$
@@
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. If $m(\rho)$ can be attained, then
@@
Let $\{(u_{n},v_{n})\}\subset S_{\rho}$ be a minimizing sequence for $m(\rho)$, while it does not satisfy (i). Thus,
$0<L:=\limsup_{n\rightarrow\infty}\sup_{y\in\mathbb{R}^{N}}\int_{B_{1}(y)}\left(|u_{n}|^{2}+|v_{n}|^{2}\right)\mathrm{d}x\leq\rho^{2},$ (3.10)
@@
$\left(u_{n}(\cdot-y_{n}),v_{n}(\cdot-y_{n})\right)\rightarrow$ $(u,v)\quad\text{ a.e.\leavevmode\nobreak\ in }\mathbb{R}^{N}\times\mathbb{R}^{N},$
@@
$\widetilde{u}_{n}:=u_{n}\left(\cdot-y_{n}\right)-u,\quad\widetilde{v}_{n}:=v_{n}\left(\cdot-y_{n}\right)-v.$
@@
So, by (3.12), $\left(u_{n}\left(\cdot-y_{n}\right),v_{n}\left(\cdot-y_{n}\right)\right)\rightarrow(u,v)$ in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ as $n\rightarrow\infty$. ∎
@@
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. For $\rho>0$, let $m(\rho)$ and $m^{J}(\rho)$ be given respectively in (1.7) and (1.8). Then $m(\rho)$ can be achieved if $m(\rho)<m^{J}(\rho)$.
@@
Suppose first that $m(\rho)<m^{J}(\rho)$. Let $\{(u_{n},v_{n})\}$ be a minimizing sequence of $m(\rho)$. If the sequence $\{(u_{n},v_{n})\}$ vanishes, according to Lions’ lemma [12, Lemma I.1], we have
$\left(u_{n},v_{n}\right)\rightharpoonup$ (0,0) in $(0,0)\ \text{ in }H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}),$
$\left(u_{n},v_{n}\right)\rightarrow$ (0,0) in $(0,0)\ \text{ in }L^{p}\left(\mathbb{R}^{N}\right)\times L^{p}(\mathbb{R}^{N}),\quad\text{ for }2<p<2^{**}.$
@@ 3.1 Comparison between $m(\rho)$ and $m^{J}(\rho)$
$Q(u_{t},v_{t})=\frac{\mathcal{D}_{1}\left(1+\frac{r_{2}}{r_{1}}\right)\|\Delta U_{t}\|_{2}^{r\gamma_{r}}}{\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{r_{2}}{r_{1}}\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\frac{r_{2}}{r_{1}}\|U_{t}\|_{2}^{2}}.$
By scaling, it is easy to get $Q(u_{t},v_{t})>\mathcal{D}_{1}$ as $t$ is large enough. The proof is done. ∎
@@
    When $R<\infty$, there hold that $m(\rho)=m^{J}(\rho)$ if $\rho\in(0,\rho^{*}]$, and $m(\rho)<m^{J}(\rho)$ if $\rho>\rho^{*}$. Moreover, when $\rho\in(0,\rho^{*})$, $m(\rho)$ is never achieved.
@@
    When $R=\infty$, we have $m(\rho)<m^{J}(\rho)$ for all $\rho>0$.
@@
On the other hand, for $10<t<$ 1, we have
@@
$-\frac{\alpha_{1}^{2}\rho^{2}}{8}=m(\rho)=I(u,v)=-\frac{\alpha_{1}^{2}\rho^{2}}{8}+\left(\frac{r}{2}-1\right)\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
@@
$\rho^{*}=\begin{cases}\left(\frac{1}{2\beta R}\right)^{\frac{1}{r-2}}\quad&\mbox{if}\;R<\infty;\\ 0\quad&\mbox{if}\;R=\infty.\end{cases}$
@@
$\rho\in\begin{cases}(\rho^{*},\infty)\quad&\mbox{if}\;\;r<\bar{r};\\ \left(\rho^{*},\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r}.\end{cases}$
Using Proposition 3.6, $m(\rho)$ can be achieved. By Proposition 3.8 again, when $\rho<\rho^{*}$, $m(\rho)$ is never achieved. Therefore we can complete the proof. ∎
@@ 3.2 Estimates on $R$
According to Theorem 1.1, $m(\rho)$ can be achieved for all $\rho>0$ if $\rho^{*}=0$. Thus, it is very important for us to know what conditions can guarantee $\rho^{*}=0$ or $\rho^{*}>0$. For that, we will discuss the value of $R$ by two cases: $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$.
@@
$\sup_{w\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|w\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha}{2}\right)w\right\|_{2}^{2}\|w\|_{2}^{r-2}}=\left(\frac{2}{\alpha}\right)^{2-\frac{(r-2)N}{4}}\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+1\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}.$ (3.29)
@@ 4 Mass-supercritical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r_{1}+r_{2}\in(\bar{r},2^{**})$. We will work in the radial Sobolev space
$H^{2}_{r}(\mathbb{R}^{N}):=\left\{u\in H^{2}(\mathbb{R}^{N}):u\mbox{ is radially symmetric}\right\}$
endowed with the equivalent norm $\|u\|:=\left(\|\Delta u\|^{2}_{2}+\|u\|^{2}_{2}\right)^{\frac{1}{2}}$. For the sake of the compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ with $2<p<2^{**}$, we assume $N\geq 2$. Furthermore, we denote
@@
=:absent:=:= : $h\left((\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2})^{\frac{1}{2}}\right),$
@@
$h(t):=\frac{1}{2}t^{2}-\frac{1}{2}\mathcal{D}_{2}\rho t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}}.$
To understand the geometry structure of the functional $I|_{S_{\rho}}$, we analyze the function $h$.
@@
$\phi(t):=\frac{1}{2}t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}-1},$
@@
$h(\bar{t})>0\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \phi(\bar{t})>\frac{1}{2}\mathcal{D}_{2}\rho\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \beta\rho^{r-2}<c^{*}(N,r).$
@@
and it is clear that there exists some $(u_{0},v_{0})\in A_{\rho}$ such that $I(u_{0},v_{0})<0$. Furthermore, we can easily find another point $(u_{1},v_{1})\in S^{r}_{\rho}\backslash A_{\rho}$ such that $I(u_{1},v_{1})<0$. Now, we can define mountain pass set as
@@
$P(u,v):=2\int_{\mathbb{R}^{N}}(|\Delta u|^{2}+|\Delta v|^{2})\mathrm{d}x-\alpha_{1}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\mathrm{d}x-\alpha_{2}\int_{\mathbb{R}^{N}}|\nabla v|^{2}\mathrm{d}x-2\beta r\gamma_{r}\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
@@
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r\in(\bar{r},2^{**})$, $N\geq 2$ and $\beta\rho^{r-2}<\min\{c^{*}(N,r),c_{*}(N,r)\}$. Let $(u_{n},v_{n})\subset S^{r}_{\rho}$ be a Palais-Smale sequence for $I|_{S_{\rho}^{r}}$ at some level $c>0$ with $P(u_{n},v_{n})\rightarrow 0$. Then up to a subsequence, $(u_{n},v_{n})\rightarrow(u,v)$ in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$, and $(u,v)$ is a radial solution of problem (1.1)-(1.2) for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
$\left(u_{n},v_{n}\right)\rightarrow$ $(u,v)\quad\text{ a.e.\leavevmode\nobreak\ in }\mathbb{R}^{N}\times\mathbb{R}^{N}.$
@@
Next, we claim that $\lambda>0$, $u\neq 0$ and $v\neq 0$. In fact, by (4.5), $\lambda_{n}\rightarrow\lambda$, $P(u_{n},v_{n})\rightarrow 0$ and $r\in(\bar{r},2^{**})$, we have
$\lambda\rho^{2}=\lim_{n\rightarrow\infty}\left(\frac{\alpha_{1}}{2}\|\nabla u_{n}\|^{2}_{2}+\frac{\alpha_{2}}{2}\|\nabla v_{n}\|^{2}_{2}\right)+\frac{\beta}{2}\left(N-r\frac{N-4}{2}\right)\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x\geq 0.$ (4.6)
If $\lambda=0$, it follows from (4.6) that
$\lim_{n\rightarrow\infty}\|\Delta u_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\Delta v_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\nabla u_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\|\nabla v_{n}\|^{2}_{2}=\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x=0.$
So $I(u_{n},v_{n})\rightarrow 0$, which contradicts our assumption that $I(u_{n},v_{n})\rightarrow c>0$. Thus, there holds $\lambda>0$.
If $u=0$ or $v=0$, we deduce from (4.3) that
@@
Claim. $\liminf\limits_{n\rightarrow\infty}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)$ has a positive lower bound and $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
$2\sqrt{\lambda}\leq$ $\frac{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}+\lambda\frac{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}$
@@
$\|\Delta(u_{n}-u)\|^{2}_{2},\leavevmode\nobreak\ \|\Delta(v_{n}-v)\|^{2}_{2},\leavevmode\nobreak\ \|u_{n}-u\|^{2}_{2},\leavevmode\nobreak\ \|v_{n}-v\|^{2}_{2}\rightarrow 0\quad\text{ as }n\rightarrow\infty.$
@@
Under $\beta\rho^{r-2}<c^{*}(N,r)$, let $\widetilde{M}(\rho)$ be given in (4.2). By Lemma 4.2, we obtain a Palais-Smale sequence $(u_{n},v_{n})$ for $I|_{S^{r}_{\rho}}$ at the level $\widetilde{M}(\rho)$, which satisfies $P\left(u_{n},v_{n}\right)\rightarrow 0$. Using Lemma 4.2, if in addition $\beta\rho^{r-2}<c_{*}(N,r)$, we can obtain a radial solution $(u,v)$ for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. ∎

### v2

@@ 1 Introduction
problem (1.1)-(1.2) has a nontrivial radial solution with mountain pass type for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
Here we assume $N\geq 2$, because the compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ for $2<p<2^{**}$ only hold for $N\geq 2$.
This paper is organized as follows. In Section 2, we present some useful lemmas. In Section 3, we give the proof of Theorems 1.1 and 1.2. In Section 4, we are dedicated to the proof of Theorem 1.4. In the rest of this paper, unless otherwise specified, we always assume that $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $r_{1}$, $r_{2}>1$.
@@ 2 Preliminaries
This section is devoted to some notations and preliminary results.
@@ 3 Mass-subcritical and mass-critical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r=r_{1}+r_{2}\in(2,\bar{r}]$. Initially, we show some important properties about $m(\rho)$ given in (1.8).
@@
Let $m^{J}(\rho)$ be given in (1.9), which is well-defined for all $\rho>0$. he following result gives an accurate $m^{J}(\rho)$, and it is not attainable.
@@
$\frac{\alpha_{1}^{2}}{8}\rho^{2}=J(u,v)\geq M(\rho_{1},\rho_{2})=-\frac{\alpha_{1}^{2}}{8}\rho_{1}^{2}-\frac{\alpha_{2}^{2}}{8}\rho_{2}^{2}\geq-\frac{\alpha_{1}^{2}}{8}\rho^{2},$
@@
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. Suppose that $m(\rho)$ is attained, then
@@
Let $\{(u_{n},v_{n})\}\subset S_{\rho}$ be a minimizing sequence for $m(\rho)$, which does not satisfy (i). Thus,
$0<L=\limsup_{n\rightarrow\infty}\sup_{y\in\mathbb{R}^{N}}\int_{B_{1}(y)}\left(|u_{n}|^{2}+|v_{n}|^{2}\right)\mathrm{d}x\leq\rho^{2},$ (3.10)
@@
$\left(u_{n}(\cdot-y_{n}),v_{n}(\cdot-y_{n})\right)\rightarrow$ $(u,v)\quad\text{ a.e.\penalty 10000\ in }\mathbb{R}^{N}\times\mathbb{R}^{N},$
@@
$\widetilde{u}_{n}=u_{n}\left(\cdot-y_{n}\right)-u,\quad\widetilde{v}_{n}=v_{n}\left(\cdot-y_{n}\right)-v.$
@@
So, by (3.12), $\left(u_{n}\left(\cdot-y_{n}\right),v_{n}\left(\cdot-y_{n}\right)\right)\rightarrow(u,v)$ in $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$ as $n\rightarrow\infty$. Hence $I(u,v)=m(\rho)$, and by Lemma 3.3, $(u,v)\in S_{\rho}$ is a global minimizer. ∎
@@
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. For $\rho>0$, let $m(\rho)$ and $m^{J}(\rho)$ be given respectively in (1.8) and (1.9). Then $m(\rho)$ is achieved if $m(\rho)<m^{J}(\rho)$.
@@
Suppose that $m(\rho)<m^{J}(\rho)$. Let $\{(u_{n},v_{n})\}$ be a minimizing sequence of $m(\rho)$. If the sequence $\{(u_{n},v_{n})\}$ vanishes, according to Lions’ lemma [12, Lemma I.1], we have
$\left(u_{n},v_{n}\right)\rightharpoonup$ $(0,0)\ \text{ in }H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}),$
$\left(u_{n},v_{n}\right)\rightarrow$ $(0,0)\ \text{ in }L^{p}\left(\mathbb{R}^{N}\right)\times L^{p}(\mathbb{R}^{N}),\quad\text{ for }2<p<2^{**}.$
@@ 3.1 Comparison between $m(\rho)$ and $m^{J}(\rho)$
$Q(u_{t},v_{t})=$ $\frac{\mathcal{D}_{1}\left(1+\frac{r_{2}}{r_{1}}\right)\|\Delta U_{t}\|_{2}^{r\gamma_{r}}}{\left\|\left(\Delta+\frac{\alpha_{1}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{r_{2}}{r_{1}}\left\|\left(\Delta+\frac{\alpha_{2}}{2}\right)U_{t}\right\|_{2}^{2}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\frac{r_{2}}{r_{1}}\|U_{t}\|_{2}^{2}}$
$=$ $\frac{\mathcal{D}_{1}\left(1+\frac{r_{2}}{r_{1}}\right)t^{4-N}\|\Delta U\|_{2}^{2}}{\left(1+\frac{r_{2}}{r_{1}}\right)t^{4-N}\|\Delta U\|_{2}^{2}-\left(\alpha_{1}+\frac{r_{2}}{r_{1}}\alpha_{2}\right)t^{2-N}\|\nabla U\|_{2}^{2}+\left(\frac{\alpha_{1}^{2}}{4}+\frac{r_{2}\alpha_{2}^{2}}{4r_{1}}+\frac{\alpha_{1}^{2}-\alpha_{2}^{2}}{4}\right)t^{-N}\|U\|_{2}^{2}}.$
It is easy to get $Q(u_{t},v_{t})>\mathcal{D}_{1}$ as $t$ is large enough. The proof is done. ∎
@@
Assume either $r\in(2,\bar{r})$ or $r=\bar{r}$, $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$. Then we have
@@
    when $R<\infty$, there hold that $m(\rho)=m^{J}(\rho)$ if $\rho\in(0,\rho^{*}]$, and $m(\rho)<m^{J}(\rho)$ if $\rho>\rho^{*}$. Moreover, when $\rho\in(0,\rho^{*})$, $m(\rho)$ is never achieved;
@@
    when $R=\infty$, there hold that $m(\rho)<m^{J}(\rho)$ for all $\rho>0$.
@@
On the other hand, for $0<t<1$, we have
@@
$\frac{\alpha_{1}^{2}\rho^{2}}{8}=m(\rho)=I(u,v)=-\frac{\alpha_{1}^{2}\rho^{2}}{8}+\left(\frac{r}{2}-1\right)\beta\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
@@
[…]
@@
[…]
Using Proposition 3.6, $m(\rho)$ is achieved. By Proposition 3.8 again, when $\rho<\rho^{*}$, $m(\rho)$ is never achieved. Therefore we can complete the proof. ∎
@@ 3.2 Estimates on $R$
According to Theorem 1.1, $m(\rho)$ is achieved for all $\rho>0$ if $\rho^{*}=0$. Thus, it is very important for us to know what conditions can guarantee $\rho^{*}=0$ or $\rho^{*}>0$. For that, we will discuss the value of $R$ by two cases: $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$.
@@
$sup_{w\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|w\|_{r}^{r}}{\left\|\left(\Delta+\frac{\alpha}{2}\right)w\right\|_{2}^{2}\|w\|_{2}^{r-2}}=\left(\frac{2}{\alpha}\right)^{2-\frac{(r-2)N}{4}}\sup_{u\in H^{2}(\mathbb{R}^{N})\backslash\{0\}}\frac{\|u\|_{r}^{r}}{\left\|\left(\Delta+1\right)u\right\|_{2}^{2}\|u\|_{2}^{r-2}}.$ (3.29)
@@ 4 Mass-supercritical case
In this section, we consider the existence of normalized solution for problem (1.1)-(1.2) with $r_{1}+r_{2}\in(\bar{r},2^{**})$. We will work in the space $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$ as in (1.12). For the sake of the compact embedding $H_{r}^{2}(\mathbb{R}^{N})\subset L^{p}(\mathbb{R}^{N})$ with $2<p<2^{**}$, we will assume $N\geq 2$. Furthermore, we denote
@@
$=:$ $h\left((\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2})^{\frac{1}{2}}\right),$
@@
$h(t)=\frac{1}{2}t^{2}-\frac{1}{2}\mathcal{D}_{2}\rho t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}}.$
To understand the geometry structure of the functional $I|_{S_{\rho}}$, we shall analyze the function $h$.
@@
$\phi(t)=\frac{1}{2}t-\mathcal{D}_{1}\beta\rho^{r\left(1-\gamma_{r}\right)}t^{r\gamma_{r}-1},$
@@
$h(\bar{t})>0\penalty 10000\ \Leftrightarrow\penalty 10000\ \phi(\bar{t})>\frac{1}{2}\mathcal{D}_{2}\rho\penalty 10000\ \Leftrightarrow\penalty 10000\ \beta\rho^{r-2}<c^{*}(N,r).$
@@
[…]
@@
$P(u,v)=2\int_{\mathbb{R}^{N}}(|\Delta u|^{2}+|\Delta v|^{2})\mathrm{d}x-\alpha_{1}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\mathrm{d}x-\alpha_{2}\int_{\mathbb{R}^{N}}|\nabla v|^{2}\mathrm{d}x-2\beta r\gamma_{r}\int_{\mathbb{R}^{N}}|u|^{r_{1}}|v|^{r_{2}}\mathrm{d}x.$
@@
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r\in(\bar{r},2^{**})$, $N\geq 2$ and $\beta\rho^{r-2}<\min\{c^{*}(N,r),c_{*}(N,r)\}$. Let $(u_{n},v_{n})\subset S^{r}_{\rho}$ be a Palais-Smale sequence of $I|_{S_{\rho}^{r}}$ at some level $c>0$ with $P(u_{n},v_{n})\rightarrow 0$. Then up to a subsequence, $(u_{n},v_{n})\rightarrow(u,v)$ in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$, and $(u,v)$ is a radial solution of problem (1.1)-(1.2) for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
$\left(u_{n},v_{n}\right)\rightarrow$ $(u,v)\quad\text{ a.e.\penalty 10000\ in }\mathbb{R}^{N}\times\mathbb{R}^{N}.$
@@
Next, we claim that $u\neq 0$ and $v\neq 0$. If $u=0$ or $v=0$, we deduce from (4.3) that
@@
Claim. $\liminf\limits_{n\rightarrow\infty}\left(\|\Delta u_{n}\|^{2}_{2}+\|\Delta v_{n}\|^{2}_{2}\right)$ is positive and $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$.
@@
$2\sqrt{\lambda}\leq$ $\frac{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}+\lambda\frac{(\|u_{n}-u\|^{2}_{2}+\|v_{n}-v\|^{2}_{2})^{\frac{1}{2}}}{\left(\|\Delta(u_{n}-u)\|^{2}_{2}+\|\Delta(v_{n}-v)\|^{2}_{2}\right)^{\frac{1}{2}}}$ (4.12)
@@
$\|\Delta(u_{n}-u)\|^{2}_{2},\penalty 10000\ \|\Delta(v_{n}-v)\|^{2}_{2},\penalty 10000\ \|u_{n}-u\|^{2}_{2},\penalty 10000\ \|v_{n}-v\|^{2}_{2}\rightarrow 0\quad\text{ as }n\rightarrow\infty.$
@@
Under $\beta\rho^{r-2}<c^{*}(N,r)$, let $\widetilde{M}(\rho)$ be given in (4.2), which is positive. By Lemma 4.2, we obtain a Palais-Smale sequence $(u_{n},v_{n})$ of $I|_{S^{r}_{\rho}}$ at the level $\widetilde{M}(\rho)$, which satisfies $P\left(u_{n},v_{n}\right)\rightarrow 0$. Using Lemma 4.3, if in addition $\beta\rho^{r-2}<c_{*}(N,r)$, we obtain a radial solution $(u,v)$ for some $\lambda>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. ∎
Acknowledgements. Z. Jin was supported by Fundamental Research Program of Shanxi Province of China(No. 202303021212160) and the China Scholarship Council(No. 202508140048), G. Wang was supported by Fundamental Research Program of Shanxi Province of China (No. 202403021221163). Last we thank the anonymous reviewer for the careful reading of our manuscript and the insightful comments and suggestions.